15% of the human body mass is bone and 65% of the mass of human bone is Calcium. How many moles of calcium are there in a 155 pound person?


1
Expert's answer
2021-10-08T02:10:07-0400

We use the information given and with unit factors, we can find the amount of Ca expressed in mol:


"155\text{ pounds} \times \cfrac{15\text{ pounds of bone}}{100\text{ pounds of human body}} \times \cfrac{65\text{ pounds of Ca}}{100\text{ pounds of human bone}} \times \cfrac{453.592\text{ g Ca}}{1\text{ pound Ca}} \times \cfrac{1\text{ mol Ca}}{40.078\text{ g Ca}}=171.04\text{ mol Ca}"


In conclusion, 155 pounds of a human body have around 171.04 mol of Ca.

Reference

  • Chang, R., & Goldsby, K. A. (2010). Chemistry. Chemistry, 10th ed.; McGraw-Hill Education: New York, NY, USA.